# PoisonedRAG: Knowledge Poisoning Attacks to Retrieval-Augmented Generation of Large Language Models

### v1

@@ 5.1 Experimental Setup
Datasets: We use three benchmark question-answering datasets in our evaluation: Natural Questions (NQ) [36], HotpotQA [37], and MS-MARCO [38], where each dataset has a knowledge database. The knowledge databases of NQ and HotpotQA are collected from Wikipedia, which contains 2,681,468 and 5,233,329 texts, respectively. The knowledge database of MS-MARCO is collected from web documents using the MicroSoft Bing search engine [80], which contains 8,841,823 texts. Each dataset also contains a set of questions. Table 12 (in Appendix) shows statistics of datasets.
@@
RAG Setup:  Recall the three components in RAG: knowledge database, retriever, and LLM. Their setups are as below:
@@
  • Knowledge database: We use the knowledge database of each dataset as that for RAG, i.e., we have 3 knowledge databases in total.
@@
  • Retriever: We consider three retrievers: Contriever [30], Contriever-ms (fine-tuned on MS-MARCO) [30], and ANCE [31]. Following previous studies [41, 14], by default, we use the dot product between the embedding vectors of a question and a text in the knowledge database to calculate their similarity score. We will also study the impact of this factor in our evaluation.
@@
  • LLM:  We consider PaLM 2 [3], GPT-4 [2], GPT-3.5-Turbo [1], LLaMA-2 [39] and Vicuna [81]. The system prompt used to let an LLM generate an answer for a question can be found in Appendix B. We set the temperature parameter of LLM to be 0.1.
@@
Unless otherwise mentioned, we adopt the following default setting. We use the NQ knowledge database and the Contriever retriever. Following previous study [14], we retrieve $5$ most similar texts from the knowledge database as the context for a question. Moreover, we calculate the dot product between the embedding vectors of a question and each text in the knowledge database to measure their similarity. We use PaLM 2 as the default LLM as it is very powerful (with 540B parameters) and free of charge, enabling us to conduct systematic evaluations. We will evaluate the impact of each factor on our knowledge poisoning attacks.
@@
Target questions and answers:  PoisonedRAG aims to make RAG produce attacker-chosen target answers for attacker-chosen target questions. Following the evaluation of previous studies [69, 82, 83, 84] on targeted poisoning attacks, we randomly select some target questions in each experiment trial and repeat the experiment multiple times. In particular, we randomly select $10$ close-ended questions from each dataset as the target questions. Moreover, we repeat the experiments 10 times (we exclude questions that are already selected when repeating the experiment), resulting in 100 target questions in total. We select close-ended questions (e.g., “Who is the CEO of OpenAI?") rather than open-ended questions (we defer the discussion on open-ended questions to Section 7) because we aim to quantitatively evaluate the effectiveness of our attacks since close-ended questions have specific, factual answers. In Appendix A, we show a set of selected target questions. For each target question, we use GPT-4 to randomly generate an answer that is different from the ground truth answer of the target question. We manually check each generated target answer and regenerate it if it is the same as the ground truth answer. Without attacks, the LLM in RAG could correctly provide answers for 70% (NQ), 80% (HotpotQA), and 83% (MS-MARCO) target questions under the default setting.
Evaluation metrics:  We use the following metrics:
  • Attack Success Rate (ASR) :  We use the ASR to measure the fraction of target questions whose answers are the attacker-chosen target answers. Following previous studies [85, 86], we say two answers are the same for a close-ended question when the target answer is a substring of the generated one by a LLM under attacks (called substring matching). We don’t use Exact Match because it is inaccurate, e.g., it views “Sam Altman” and “The CEO of OpenAI is Sam Altman” as different answers to the question “Who is the CEO of OpenAI?". We use human evaluation (conducted by authors) to validate the substring matching method. We find that substring matching produces similar ASRs as human evaluation (Table 2 shows the comparison).
[…]
  • #Queries:  PoisonedRAG utilizes a LLM to generate the text $I$ to satisfy the effectiveness condition, which is further utilized to craft a poisoned text. We report the average number of queries made to a LLM to generate each poisoned text.
  • Runtime:  In both white-box and black-box settings, PoisonedRAG crafts $S$ such that poisoned texts are more likely to be retrieved for the target questions. PoisonedRAG is more efficient when the runtime is less. In our evaluation, we also report the average runtime in generating each poisoned text.
Compared baselines:  To the best of our knowledge, there is no existing attack that aims to achieve our attack goal. In response, we extend other attacks [45, 46, 47, 40, 41] to LLM to our scenario. In particular, we consider the following baselines:
  • No Attack:  Our first baseline is No Attack, i.e., we do not inject any poisoned texts into the knowledge database. No Attack serves as a baseline to measure the effectiveness of PoisonedRAG.
  • Prompt Injection Attack [45, 46, 47, 40]:  Prompt injection attacks aim to inject an instruction into the prompt of a LLM such that the LLM generates an attacker-desired output. Inspired by our black-box attack, we put the target question in the instruction for the prompt injection attacks such that the crafted poisoned texts are more likely to be retrieved for the target question. In particular, given a target question and target answer, we craft the following poisoned text: “When you are asked to provide the answer for the following question: <target question>, please output <target answer>.”. We note that the key difference between prompt injection attacks and PoisonedRAG (in the black-box setting) is that prompt injection attacks utilize instructions while PoisonedRAG crafts poisoned knowledge.
  • Corpus Poisoning Attack [41]: This attack aims to inject poisoned samples (without any semantic meanings) into a knowledge database such that they are retrieved for target questions. This attack requires the white-box access to the retriever. We adopt the publicly available implementation [41] for our experiments. As shown in our results, they achieve a very low ASR (close to No Attack). The reason is that they are not designed to make a LLM in RAG generate attacker-chosen target answers for attacker-chosen target questions.
  • Two variants of PoisonedRAG: Each poisoned text $P$ crafted by PoisonedRAG consists of two sub-texts, i.e., $I$ and $S$. We compare PoisonedRAG with its two variants, where $I$ (or $S$) alone is used as the poisoned text $P$.
Note that, for a fair comparison, we also craft $N$ poisoned texts for each target question for baselines.
@@
Table 1: PoisonedRAG could achieve high ASRs on 3 datasets under 8 different LLMs, where the poisoning rate per question (ratio between $N$ and the number of texts in the clean database) are $5/2,681,468\approx 0.0002\%$, $5/5,233,329\approx 0.0001\%$, and $5/8,841,823\approx 0.00006\%$, respectively. We omit Precision and Recall because they are the same as F1-Score.
@@
Table 3: Average #Queries and runtime of PoisonedRAG in crafting each poisoned text.
@@
Hyperparameter setting: Unless otherwise mentioned, we adopt the following hyperparameters for PoisonedRAG. We inject $N=5$ poisoned texts for each target question. Recall that, in both black-box and white-box attacks, we use a LLM to generate $I$. We use GPT-4 in our experiment, where the temperature parameter is set to be 1. Moreover, we set the maximum number of trials $L=50$ when using LLM to generate $I$. We set the length of $I$ to be $V=30$. In our white-box attack, we use HotFlip [75], a state-of-the-art method to craft adversarial texts, to solve the optimization problem in Equation 6. We will conduct a systematic evaluation on the impact of these hyperparameters on PoisonedRAG.
@@ 5.2 Main Results
PoisonedRAG achieves high ASRs and F1-Score: Table 1 shows the ASRs of PoisonedRAG under black-box and white-box settings. We have the following observations from the experimental results. First, PoisonedRAG could achieve high ASRs on different datasets and LLMs under both white-box and black-box settings when injecting 5 poisoned texts for each target question into a knowledge database with millions of texts. For instance, in the black-box setting, PoisonedRAG could achieve 97% (on NQ), 99% (on HotpotQA), and 91% (on MS-MARCO) ASRs for RAG with PaLM 2. Our experimental results demonstrate that RAG is extremely vulnerable to our knowledge poisoning attacks. Second, PoisonedRAG achieves high F1-Scores under different settings, e.g., larger than 90% in almost all cases. The results demonstrate that the poisoned texts crafted by PoisonedRAG are very likely to be retrieved for target questions, which is also the reason why PoisonedRAG could achieve high ASRs. Third, in most cases, PoisonedRAG is more effective in the white-box setting compared to the black-box setting. This is because PoisonedRAG can leverage more knowledge of the retriever in the white-box setting, and hence the crafted poisoned text has a larger similarity with a target question and is more likely to be retrieved, e.g., the F1-Score of the PoisonedRAG under the white-box setting is higher than that of the black-box setting. We note that PoisonedRAG achieves better ASRs in the black-box setting than the white-box setting in some cases. We suspect the reason is that HotFlip (used to craft adversarial texts in the white-box setting) slightly influences the semantics of poisoned texts in these cases.
Our substring matching metric achieves similar ASRs to human evaluation: We use substring matching to calculate ASR in our evaluation. We conduct a human evaluation to validate such a method, where we manually check whether a LLM in RAG produces the attacker-chosen target answer for each target question. Table 2 shows the results. We find that ASR calculated by substring matching is similar to that of human evaluation, demonstrating the reliability of the substring matching evaluation metric. We note that it is still an open challenge to develop a perfect metric.
PoisonedRAG is computationally efficient:  Table 3 shows the average #Queries and runtime of PoisonedRAG. We have two key observations. First, on average, PoisonedRAG only needs to make around 2 queries to the GPT-4 LLM to craft each poisoned text. Second, it takes far less than 1 second for PoisonedRAG to optimize the poisoned text such that it has a large semantic similarity with the target question in the black-box setting. The reason is that PoisonedRAG directly concatenates the text generated by a LLM and the target question to craft a poisoned text, which is very efficient. Further, it takes less than 30 seconds to optimize each poisoned text in the white-box setting. We note that PoisonedRAG could craft poisoned texts in parallel.
@@
NQ No Attack 0.01 NA
@@
HotpotQA No Attack 0.01 NA
@@
MS-MARCO No Attack 0.03 NA
@@
Table 5: PoisonedRAG outperforms its two variants.
Dataset Attack $S\oplus I$ $S$ $I$
ASR F1-Score ASR F1-Score ASR F1-Score
NQ PoisonedRAG (Black-Box) 0.97 0.96 0.03 1.0 0.69 0.48
PoisonedRAG (White-Box) 0.97 1.0 0.02 0.99 0.51 0.93
Hotpot QA PoisonedRAG (Black-Box) 0.99 1.0 0.06 1.0 1.0 0.99
PoisonedRAG (White-Box) 0.94 1.0 0.08 1.0 0.71 0.99
MS-MA RCO PoisonedRAG (Black-Box) 0.91 0.89 0.02 1.0 0.57 0.36
PoisonedRAG (White-Box) 0.90 0.94 0.06 0.97 0.47 0.87
PoisonedRAG outperforms baselines: Table 4 compares PoisonedRAG with baselines under the defaualt setting. We have the following observations. First, the Corpus Poisoning Attack achieves a very low ASR (close to No Attack) because it is not designed to make a LLM in RAG produce attacker-chosen target questions for attacker-chosen target answers. Note that Corpus Poisoning Attack is similar to PoisonedRAG (white-box setting) when PoisonedRAG uses $S$ alone as the poisoned text $P$ (i.e., $P=S$). Second, prompt injection attack also achieves a non-trivial ASR, although it is worse than PoisonedRAG. The reason is that, inspired by PoisonedRAG in the black-box setting, we also add the target question to the poisoned texts crafted by prompt injection attacks. As a result, some poisoned texts crafted by prompt injection attacks could be retrieved for the target questions as reflected by a non-trivial F1-Score. As LLMs are good at following instructions, prompt injection attack achieves a non-trivial ASR. Note that the key difference between PoisonedRAG and prompt injection attack is that PoisonedRAG relies on malicious knowledge instead of instructions to mislead LLMs. Third, PoisonedRAG achieves better ASRs than baselines, demonstrating the effectiveness of PoisonedRAG.
PoisonedRAG outperforms its two variants:  Table 5 shows the comparison results of PoisonedRAG with its two variants, where we use $I$ (or $S$) alone as the poisoned text, respectively. The results show PoisonedRAG achieves a higher ASR than its two variants, demonstrating that both $I$ and $S$ are needed by PoisonedRAG to be effective. Note that, when $S$ alone is used as the poisoned text, the F1-Score is very high but ASR is very low. The reason is that $S$ has a high similarity score with a target question but cannot make a LLM generate the target answer when used as the context.
5.3 Ablation Study
In this section, we study the impact of hyperparameters on our PoisonedRAG. Due to limited space, we defer the results with different LLMs used in RAG to Appendix D.
5.3.1 Impact of Hyperparameters in RAG
Impact of retriever:  Table 6 shows the effectiveness of PoisonedRAG for different retrievers under the default setting. Our results demonstrate that PoisonedRAG is consistently effective for different retrievers. PoisonedRAG is consistently effective in the black-box setting because the crafted poisoned texts are semantically similar to the target questions. Thus, they are very likely to be retrieved for the target questions by different retrievers, e.g., F1-Score is consistently high.
Impact of $k$:  Figure 3 shows the impact of $k$. We have following observations. First, ASR of PoisonedRAG is consistently high when $k\leq N$ ($N=5$ by default). The reason is that most of the retrieved texts are poisoned ones when $k\leq N$, e.g., Precision (measure the fraction of retrieved texts that are poisoned ones) is very high and Recall increases as $k$ increases. When $k>N$, ASR (or Precision) decreases as $k$ increases. The reason is that $(k-N)$ retrieved texts are clean ones as the total number of poisoned texts for each target question is $N$. Note that Recall is close to $1$ when $k>N$, which means almost all poisoned texts are retrieved for target questions.

### v2

@@ 5.1 Experimental Setup
Datasets. We use three benchmark question-answering datasets in our evaluation: Natural Questions (NQ) [38], HotpotQA [39], and MS-MARCO [40], where each dataset has a knowledge database. The knowledge databases of NQ and HotpotQA are collected from Wikipedia, which contains 2,681,468 and 5,233,329 texts, respectively. The knowledge database of MS-MARCO is collected from web documents using the MicroSoft Bing search engine [84], which contains 8,841,823 texts. Each dataset also contains a set of questions. Table 14 (in Appendix) shows statistics of datasets.
@@
RAG Setup.  Recall the three components in RAG: knowledge database, retriever, and LLM. Their setups are as below:
@@
  • Knowledge database. We use the knowledge database of each dataset as that for RAG, i.e., we have 3 knowledge databases in total.
@@
  • Retriever. We consider three retrievers: Contriever [32], Contriever-ms (fine-tuned on MS-MARCO) [32], and ANCE [33]. Following previous studies [43, 14], by default, we use the dot product between the embedding vectors of a question and a text in the knowledge database to calculate their similarity score. We will also study the impact of this factor in our evaluation.
@@
  • LLM.  We consider PaLM 2 [3], GPT-4 [2], GPT-3.5-Turbo [1], LLaMA-2 [41] and Vicuna [85]. The system prompt used to let an LLM generate an answer for a question can be found in Appendix B. We set the temperature parameter of LLM to be 0.1.
@@
Unless otherwise mentioned, we adopt the following default setting. We use the NQ knowledge database and the Contriever retriever. Following previous study [14], we retrieve $5$ most similar texts from the knowledge database as the context for a question. Moreover, we calculate the dot product between the embedding vectors of a question and each text in the knowledge database to measure their similarity. We use PaLM 2 as the default LLM as it is very powerful (with 540B parameters) and free of charge, enabling us to conduct systematic evaluations. We will evaluate the impact of each factor on our knowledge corruption attacks.
@@
Target questions and answers.  PoisonedRAG aims to make RAG produce attacker-chosen target answers for attacker-chosen target questions. Following the evaluation of previous studies [70, 86, 87, 88] on targeted poisoning attacks, we randomly select some target questions in each experiment trial and repeat the experiment multiple times. In particular, we randomly select $10$ close-ended questions from each dataset as the target questions. Moreover, we repeat the experiments 10 times (we exclude questions that are already selected when repeating the experiment), resulting in 100 target questions in total. We select close-ended questions (e.g., “Who is the CEO of OpenAI?") rather than open-ended questions (we defer the discussion on open-ended questions to Section 8) because we aim to quantitatively evaluate the effectiveness of our attacks since close-ended questions have specific, factual answers. In Appendix A, we show a set of selected target questions. For each target question, we use GPT-4 to randomly generate an answer that is different from the ground truth answer of the target question. We manually check each generated target answer and regenerate it if it is the same as the ground truth answer. Without attacks, the LLM in RAG could correctly provide answers for 70% (NQ), 80% (HotpotQA), and 83% (MS-MARCO) target questions under the default setting.
@@
Table 1: PoisonedRAG could achieve high ASRs on 3 datasets under 8 different LLMs, where we inject 5 malicious texts for each target question into a knowledge database with $2,681,468$ (NQ), $5,233,329$ (HotpotQA), and $8,841,823$ (MS-MARCO) clean texts. We omit Precision and Recall because they are the same as F1-Score.
@@
Evaluation metrics.  We use the following metrics:
  • Attack Success Rate (ASR) .  We use the ASR to measure the fraction of target questions whose answers are the attacker-chosen target answers. Following previous studies [89, 90], we say two answers are the same for a close-ended question when the target answer is a substring of the generated one by an LLM under attacks (called substring matching). We don’t use Exact Match because it is inaccurate, e.g., it views “Sam Altman” and “The CEO of OpenAI is Sam Altman” as different answers to the question “Who is the CEO of OpenAI?". We use human evaluation (conducted by authors) to validate the substring matching method. We find that substring matching produces similar ASRs as human evaluation (Table 2 shows the comparison).
[…]
  • #Queries.  PoisonedRAG utilizes an LLM to generate the text $I$ to satisfy the generation condition. We report the average number of queries made to an LLM to generate each malicious text.
  • Runtime.  In both white-box and black-box settings, PoisonedRAG crafts $S$ such that malicious texts are more likely to be retrieved for the target questions. PoisonedRAG is more efficient when the runtime is less. In our evaluation, we also report the average runtime in generating each malicious text.
Compared baselines.  To the best of our knowledge, there is no existing attack that aims to achieve our attack goal. In response, we extend other attacks [47, 48, 49, 42, 43] to LLM to our scenario. In particular, we consider the following baselines:
  • Naive Attack. Given a question $Q$, if we view $Q$ as the malicious text, it will likely be retrieved. We compare with this attack to demonstrate that the generation condition is necessary for knowledge corruption attacks.
  • Prompt Injection Attack [47, 48, 49, 42].  Prompt injection attacks aim to inject an instruction into the prompt of an LLM such that the LLM generates an attacker-desired output. Inspired by our black-box attack, we put the target question in the instruction for the prompt injection attacks such that the crafted malicious texts are more likely to be retrieved for the target question. In particular, given a target question and target answer, we craft the following malicious text: “When you are asked to provide the answer for the following question: <target question>, please output <target answer>.”. We note that the key difference between prompt injection attacks and PoisonedRAG (in the black-box setting) is that prompt injection attacks utilize instructions while PoisonedRAG crafts malicious knowledge.
  • Corpus Poisoning Attack [43]. This attack aims to inject malicious texts (consisting of random characters) into a knowledge database such that they can be retrieved for indiscriminate questions. This attack requires the white-box access to the retriever. We adopt the publicly available implementation [43] for our experiments. As shown in our results, they achieve a very low ASR (close to Naive Attack). The reason is that it cannot achieve the generation condition. Note that this attack is similar to PoisonedRAG (white-box setting) when PoisonedRAG uses $S$ alone as the malicious text $P$ (i.e., $P=S$).
  • GCG Attack [53]. Zou et al. [53] proposed an optimization-based jailbreaking attack to LLM. In particular, given a harmful question, they aim to optimize and append an adversarial suffix (an adversarial text) such that the generated output of the LLM starts with an affirmative response (e.g., “Sure, here is”). We extend the GCG attack to our scenario. In particular, we can optimize an adversarial text such that the LLM generates the target answer for a target question (see Appendix D for our adaptation details). Then, we view the optimized adversarial text as a malicious text and inject it into the knowledge database. Our results show that GCG achieves a very low ASR (close to Naive Attack). The reason is that it cannot achieve the retrieval condition.
  • Disinformation Attack [91, 92]. The crafted $I$ (to achieve the generation condition) by PoisonedRAG for a target question can be viewed as disinformation [91, 92]. Thus, we compare with this baseline where we view the crafted $I$ as a malicious text, i.e., $P=I$. This baseline can be viewed as a variant of PoisonedRAG.
Note that, for a fair comparison, we also craft $N$ malicious texts for each target question for baselines. Existing baselines are not designed to simultaneously achieve retrieval and generation conditions, resulting in sub-optimal performance.
Table 3: Average #Queries and runtime of PoisonedRAG in crafting each malicious text.
@@
Hyperparameter setting. Unless otherwise mentioned, we adopt the following hyperparameters for PoisonedRAG. We inject $N=5$ malicious texts for each target question. Recall that, in both black-box and white-box attacks, we use an LLM to generate $I$. We use GPT-4 in our experiment, where the temperature parameter is set to be 1. Moreover, we set the maximum number of trials $L=50$ when using LLM to generate $I$. We set the length of $I$ to be $V=30$. In our white-box attack, we use HotFlip [78], a state-of-the-art method to craft adversarial texts, to solve the optimization problem in Equation 5. We will conduct a systematic evaluation on the impact of these hyperparameters on PoisonedRAG.
@@ 5.2 Main Results
PoisonedRAG achieves high ASRs and F1-Score. Table 1 shows the ASRs of PoisonedRAG under black-box and white-box settings. We have the following observations from the experimental results. First, PoisonedRAG could achieve high ASRs on different datasets and LLMs under both white-box and black-box settings when injecting 5 malicious texts for each target question into a knowledge database with millions of texts. For instance, in the black-box setting, PoisonedRAG could achieve 97% (on NQ), 99% (on HotpotQA), and 91% (on MS-MARCO) ASRs for RAG with PaLM 2. Our experimental results demonstrate that RAG is extremely vulnerable to our knowledge corruption attacks. Second, PoisonedRAG achieves high F1-Scores under different settings, e.g., larger than 90% in almost all cases. The results demonstrate that the malicious texts crafted by PoisonedRAG are very likely to be retrieved for target questions, which is also the reason why PoisonedRAG could achieve high ASRs. Third, in most cases, PoisonedRAG is more effective in the white-box setting compared to the black-box setting. This is because PoisonedRAG can leverage more knowledge of the retriever in the white-box setting, and hence the crafted malicious text has a larger similarity with a target question and is more likely to be retrieved, e.g., the F1-Score of the PoisonedRAG under the white-box setting is higher than that of the black-box setting. We note that PoisonedRAG achieves better ASRs in the black-box setting than the white-box setting in some cases. We suspect there are two reasons. First, HotFlip (used to craft adversarial texts in the white-box setting) slightly influences the semantics of malicious texts in these cases. Second, prepending a target question could also contribute to the generation condition, making the black-box attack more effective when most of the malicious texts are retrieved (i.e., F1-Score is high).
Our substring matching metric achieves similar ASRs to human evaluation. We use substring matching to calculate ASR in our evaluation. We conduct a human evaluation to validate such a method, where we manually check whether an LLM in RAG produces the attacker-chosen target answer for each target question. Table 2 shows the results. We find that ASR calculated by substring matching is similar to that of human evaluation, demonstrating the reliability of the substring matching evaluation metric. We note that it is still an open challenge to develop a perfect metric.
@@
NQ Naive Attack 0.03 1.0
Corpus Poisoning Attack 0.01 0.99
Disinformation Attack 0.69 0.48
@@
HotpotQA Naive Attack 0.06 1.0
Corpus Poisoning Attack 0.01 1.0
Disinformation Attack 1.0 0.99
@@
MS-MARCO Naive Attack 0.02 1.0
Corpus Poisoning Attack 0.03 0.97
Disinformation Attack 0.57 0.36
@@
PoisonedRAG is computationally efficient.  Table 3 shows the average #Queries and runtime of PoisonedRAG. We have two key observations. First, on average, PoisonedRAG only needs to make around 2 queries to the GPT-4 to craft each malicious text. Second, it takes far less than 1 second for PoisonedRAG to optimize the malicious text in the black-box setting. The reason is that PoisonedRAG directly concatenates the text generated by an LLM and the target question to craft a malicious text. Further, it takes less than 30 seconds to optimize each malicious text in the white-box setting. We note that PoisonedRAG could craft malicious texts in parallel.
@@
PoisonedRAG outperforms baselines. Table 4 compares PoisonedRAG with baselines under the default setting. We have the following observations. First, PoisonedRAG outperforms those baselines, demonstrating the effectiveness of PoisonedRAG. The reason is that those baselines are not designed to simultaneously achieve retrieval and generation conditions. Second, prompt injection attack also achieves a non-trivial ASR, although it is worse than PoisonedRAG. The reason is that, inspired by PoisonedRAG in the black-box setting, we also add the target question to the malicious texts crafted by prompt injection attacks. As a result, some malicious texts crafted by prompt injection attacks could be retrieved for the target questions as reflected by a non-trivial F1-Score. As LLMs are good at following instructions, prompt injection attack achieves a non-trivial ASR. Note that the key difference between PoisonedRAG and prompt injection attack is that PoisonedRAG relies on malicious knowledge instead of instructions to mislead LLMs. Third, the disinformation attack (a variant of PoisonedRAG) also achieves a non-trivial ASR as some crafted malicious texts by this attack can also be retrieved (reflected by a non-trivial F1-Score). The reason is that those malicious texts are relevant to the target question. Fourth, Naive Attack, Corpus Poisoning Attack, and GCG Attack are ineffective because they cannot achieve generation, generation, and retrieval condition, respectively.
5.3 Ablation Study
We study the impact of hyperparameters on PoisonedRAG. For space reasons, we defer the results for different LLMs used in RAG to Appendix G.
5.3.1 Impact of Hyperparameters in RAG
Impact of retriever. Table 5 shows the effectiveness of PoisonedRAG for different retrievers under the default setting. Our results demonstrate that PoisonedRAG is consistently effective for different retrievers. PoisonedRAG is effective in the black-box setting because the crafted malicious texts are semantically similar to the target questions. Thus, they are very likely to be retrieved for the target questions by different retrievers, e.g., F1-Score is consistently high.
Impact of $k$.  Figure 3 shows the impact of $k$. We have the following observations. First, ASR of PoisonedRAG is high when $k\leq N$ ($N=5$ by default). The reason is that most of the retrieved texts are malicious ones when $k\leq N$, e.g., Precision (measure the fraction of retrieved texts that are malicious ones) is very high and Recall increases as $k$ increases. When $k>N$, ASR (or Precision) decreases as $k$ increases. The reason is that $(k-N)$ retrieved texts are clean ones as the total number of malicious texts for each target question is $N$. Note that Recall is close to $1$ when $k>N$, which means almost all malicious texts are retrieved for target questions.